Determine the volume occupied by 2.5 mol of a gas at 18 degrees if the pressure is at 81.8 kPa


1
Expert's answer
2022-02-24T15:49:04-0500

P = 81.8 kPa

V = unknown

n = 2.5 mol

R = gas constant = 8.314 L kPa mol-1 K-1

T = 18°C + 273.15 = 291.15 K


Solution:

The ideal gas law can be used.

PV = nRT

The ideal gas law can be rewritten as:

V = nRT / P

Plug in the known values and solve for 'V':

V = (2.5 mol × 8.314 L kPa mol-1 K-1 × 291.15 K) / (81.8 kPa) = 73.98 L ≈ 74.0 L

V = 74.0 L


Answer: 74.0 L